OpenStudy (anonymous):

Differential Equations help?? I don't see how using substitution x=lnt t u(x)=y(e^x) into t2 y'' + t y' + y = 0 we can get u''+u=0

OpenStudy (anonymous):

I was given x=lnt I don't know how they came up with u(x)=y(e^x) or what that means

OpenStudy (amistre64):

\[(t^2\ y'' + t\ y' + y = 0)/t^2\] \[y'' + t^{-1}\ y' +t^{-2} y = 0\] hmmm, whats the name of the method you think that used?

OpenStudy (amistre64):

.... think that they used. cant type

OpenStudy (amistre64):

if i were to take a guess at this, since I aint to adept at these: \[r^2+t^{-1}r+t^{-2}=0\] \[r=\frac{-t^{-1}\pm \sqrt{t^{-2}-4(t^{-2})}}{2}\] \[r=\frac{-t^{-1}\pm \sqrt{t^{-2}(1-4)}}{2}\] \[r=\frac{-t^{-1}\pm t^{-1}\sqrt{-3}}{2}\] that cant be good ....

OpenStudy (anonymous):

With substituting \(t=\ln{x}\), it follows that \(\large{\frac{dy}{dt}=\frac{dy}{dx}\frac{dx}{dt}=\frac{1}{t}\frac{dy}{dx}}\), and \(\large{\frac{d^2y}{dt^2}=\frac{1}{t^2}\frac{d^2y}{dx^2}-\frac{1}{t^2}\frac{dy}{dx}=\frac{1}{t^2}(\frac{d^2y}{dx^2}-\frac{dy}{dx})}\) (By using product rule). Substitute that into the differential equation.

OpenStudy (anonymous):

It looked like this when I was using chrome, but with firefox it's looking good!

OpenStudy (anonymous):

Let me write it again. All I did was to use the substitution \(x=\ln t\), to find the derivatives with respect to x.

OpenStudy (anonymous):

\[{dy \over dt}={1 \over t}{dy \over dx}, \text { and } {d^2y \over dt^2}={1 \over t^2}({d^2y \over dx^2}-{dy \over dx})\].

OpenStudy (anonymous):

If we substitute this into the differential equation, we get: \[{d^2y \over dx^2}-{dy \over dx}+{dy \over dx}+y=0 \implies {d^2y \over dx^2}+y=0.\] Which is the same as the equation you have.

OpenStudy (anonymous):

All you need now is to use the auxiliary equation to find the general solution, and then substitute back for \(t\).

OpenStudy (anonymous):

Sorry I'm still new at this I'm lost when you did the d^2y/dt^2=1/t^2(d^2y/dx^2−dydx)

OpenStudy (anonymous):

And when you plug into the differential equation t2 y'' + t y' + y = 0 where did the t go?

OpenStudy (anonymous):

Just use the product rule to find \(\large{\frac{d^2y}{dt^2}}\).

OpenStudy (anonymous):

The t's were cancelled out with the t's that are in the denominators of \(\large{\frac{d^2y}{dt^2}}\) and \(\large {\frac{dy}{dt}}\).

OpenStudy (anonymous):

but for d^2y/dt^2 shouldn't it be \[d^2y/dt^2=1/t^2(d^2y/dx^2 - 1/t^2 dy/dx)\]

OpenStudy (anonymous):

\[d^2y/dt^2=1/t^2(d^2y/dx^2−1/t dy/dx)\]

OpenStudy (anonymous):

sorry I meant the last one I posted shouldn't there still be a 1/t in the brackets